find the slope of the line through each pair of points, if possible, and based on the slope, indicate whether the line through the points risess from left to right, falls from left to right, is horizonal, or is vertical (-4,1)and(-3,4)and (1.5,2.6)and(0.5,(3.6)
Answer by stanbon(75887)

find the slope of the line through each pair of points, if possible, and based on the slope, indicate whether the line through the points rises from left to right, falls from left to right, is horizonal, or is vertical
(-4,1)and(-3,4)
slope = (4-1)/(-3--4) = 3/1 = 3 ;;therefore rises
------------------------

(1.5,2.6)and(0.5,(3.6)
slope = (3.6-2.6)/(0.5-1.5) = 1/-1 = -1 ;; therefore falls
--------------------------
Cheers,
Stan H.
===============